OpenStudy (anonymous):

write an equation of a line in slope intercept form that is perpendicular to y=-4x-2 and passes through the point (-16,-11)

OpenStudy (anonymous):

what do you know of perpendicular slopes? anything?

OpenStudy (anonymous):

um, i know some. but this particular type of problem gets me. im not sure to get the slope first or what. i think i have to input the pints, but when there are two sets of cordinants, i get confused. @dpaInc

OpenStudy (anonymous):

well first, the line given is perpendicular to the line you need. what is the slope of the given line? HINT: y=mx + b, "m" is the slope

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

ok... good... the relation between perpendicular slopes is that they are NEGATIVE RECIPROCALS of each other... in other words, \(\large m_1=\frac{1}{m_2} \) so if m2 is -4, then the slope perpendicular would be \(\large \frac{1}{-4} \) understand?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

good... since your question wants the equation in slope-intercept form, all we need now is the intercept because we know the slope is -1/4...

OpenStudy (anonymous):

ok that makes sense

OpenStudy (anonymous):

I see you...love the pony

OpenStudy (anonymous):

lool love the eyes

OpenStudy (anonymous):

so.... slope-intercept form: \(\large y=mx+b \) we already know that the slope is -1/4 so: \(\large y=-\frac{1}{4}x+b \) since we know the line passes through (-16, -11) plug these values into the above equation so we can solve for "b", the y-intercept...

OpenStudy (anonymous):

you should be a teacher dpalnc :)

OpenStudy (anonymous):

nah... i gave up on school... :(

OpenStudy (anonymous):

ok. -11=11/4(-16)+b

OpenStudy (anonymous):

what ?

OpenStudy (anonymous):

i think u mean this: \(\large -11=-\frac{1}{4}(-16)+b \)

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

now just solve for b

OpenStudy (anonymous):

-51?

OpenStudy (anonymous):

im only doing this so i can pass... yep... solve for "b" and put it back into the equation: \(\large y=-\frac{1}{4}x+\color{red}b \)

OpenStudy (anonymous):

-51 is not right... try again.

OpenStudy (anonymous):

-15

OpenStudy (anonymous):

there you go.. :) nice work....

OpenStudy (anonymous):

good job tyler :)

OpenStudy (anonymous):

so your equation is y = -1/4x - 15